OpenStudy (anonymous):

WILL GIVE MEDAL & FAN. What is the slope of the line given to the equation below?

OpenStudy (anonymous):

which equation missy?

OpenStudy (anonymous):

I think it's just x but idk :)

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Do you know what is the meaning of slope?

OpenStudy (anonymous):

y=mx + b ? :)

OpenStudy (anonymous):

because the general equation of a linear graph is y=mx+b m is the 'slope' and 'b' is the y-intercept

OpenStudy (anonymous):

thats not what i meant :P.. i wanted meaning of a slope.. but yes.. that equation is like a general equation of a line in slope intercept form so y = mx +b... now compare that to y =x what is m and what is b?

OpenStudy (anonymous):

Yeah so it's just 1 not x? :) @Chelsea04

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

They're intercepts ? @Mashy

OpenStudy (anonymous):

m is the slope b is the y-intercept

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

yup!